# Dynamic mapping from static labels: remote sensing dynamic sample generation with temporal-spectral embedding

### v1

@@ Abstract
Accurate remote sensing geographic mapping depends heavily on representative and timely sample data. However, rapid changes in land surface dynamics necessitate frequent updates, quickly rendering previously collected samples obsolete and imposing significant labor demands for continuous manual updates. In this study, we aim to address this problem by dynamic sample generation using existing single-date static labeled samples. We introduce TasGen, a two-stage automated framework to automatically generate dynamic samples, designed to simultaneously model spectral and temporal dependencies in time-series remote sensing imagery via temporal-spectral embedding, capturing land surface changes without additional manual annotations. Specifically, in the first stage, TasGen designs a hierarchical temporal-spectral Variational AutoEncoder (HTS-VAE) with a dual-dimension embedding to obtain two distinct stochastic latent variables that learn low-dimensional features of spectral and temporal patterns in sample data. Land surface changes show significant deviations from normal temporal-spectral patterns that often indicate environmental transitions. These changes are identified by the learned patterns from HTS-VAE in an unsupervised manner. In the second stage, TasGen trains a classifier using the stable samples with spatiotemporal features from the HTS-VAE. This trained classifier relabels the time points with detected land surface changes to generate dynamic samples applicable across multiple periods. To not only detect but also explain surface dynamics, we further propose an anomaly interpretation method based on Gibbs sampling, which attributes changes to specific spectral-temporal dimensions. We evaluate TasGen on two dynamic geographic mapping tasks (i.e., land cover mapping, wetland mapping) across six diverse global locations. TasGen achieved an average mapping OA and Kappa of 91.29% and 91.26%, considerably outperforming other state-of-the-art algorithms with a maximum of 5.92% and 6.61%. Consistent superior accuracy across all regions demonstrates TasGen’s effectiveness in tracking both long-term trends and abrupt land surface changes, while significantly reducing the need for human intervention.
@@ 1 Introduction
The Earth’s surface is undergoing unprecedented rapid changes, with urbanization encroaching upon natural habitats at an annual rate of 1.2% [Seto et al., 2012], and global wetlands decreasing by approximately 0.5% per year since the 20th century [Nicholls, 2004, Yuan et al., 2025]. Timely and accurate geographic mapping serves as a fundamental basis for monitoring ecosystem health [Soubry et al., 2021, Mariano et al., 2018], optimizing urban planning [Gong et al., 2020, Yuan et al., 2024], protecting biodiversity [Wiens et al., 2009], and understanding climate change [Yang et al., 2013]. Remote sensing has become the primary way of acquiring dynamic surface information with its capability for periodic observations [Gong, 2012, Pekel et al., 2016]. For instance, the Landsat satellite series has provided over 40 years of continuous observations, offering essential data for global land change analysis [Wulder et al., 2019].
@@
Now there are three kinds of methods addressing this challenge. The first is manually annotating multi-temporal samples. It requires visual interpretation and manual annotations at each time point, which is constrained in small study areas and short time series easily. For example, Li et al. [2017] established the first all-season sample sets (FAST) for global land cover mapping, primarily based on manual interpretation of Landsat-8 imagery regarding high-resolution Google Earth images. Yang & Huang [2021] collected training samples by integrating land cover samples derived from China’s Land-Use/Cover Datasets (CLUDs) with visually interpreted samples based on satellite time-series data, Google Earth, and Google Maps imagery. The second is using stable samples across time series to train phase-general classifiers. In other words, these methods only need manual annotations once to conduct the time-series mapping. For example, Zhang & Weng [2016] developed a method for extracting training samples from spectrally stable areas, and used Google Earth imagery as reference data for land cover identification. Huang et al. [2020] developed a novel methodology to migrate global training samples by leveraging spectral similarity and spectral distance metrics, enabling the reuse of stable training samples for dynamic global land cover mapping across multiple periods. Even though stable samples only need one-time annotation, the scarcity of stable samples and the temporal identification of stable samples limit the scaling ability. The third is the change detection method modeling the time-series patterns. Zhu & Woodcock [2014] proposed the continuous change detection and classification (CCDC) algorithm, detecting changes through unsupervised time-series modeling and training classifiers relying on static reference data from stable periods. Kennedy et al. [2010] proposed LandTrendr, considering dynamic segmentation analysis of Landsat time series and training classifiers relying on trajectory fitting and seasonal-trend decomposition. He et al. [2024] employed temporal semantic segmentation to extract land cover change information from Sentinel-2 time series data. While these change detection methods consider the temporal patterns, they struggle to adapt to serious Earth’s surface patterns because of the model initialization or dense manual annotations. Additionally, these change detection methods typically perform end-to-end detection, directly identifying change points and their types at the mapping level. However, such end-to-end frameworks often cannot guarantee sufficient accuracy due to compounded uncertainties: errors in detecting change timing or extent can directly corrupt the downstream classification. Moreover, the absence of an intermediate step on the sample level makes it difficult to evaluate and refine the model’s understanding of real surface dynamics. Without explicit sample-level transition modeling, end-to-end systems are often prone to misclassifying subtle or heterogeneous changes, especially in ecologically complex regions.
@@
Due to the complexity of both spectral and temporal dynamics inherent in time-series data, obsolete samples cannot effectively represent surface characteristics at later time points, leading to significant degradation in mapping accuracy over time. High-quality and abundant dynamic and up-to-date samples are the key to accurate time-series dynamic mapping [Stanimirova et al., 2023]. Based on the observations above, temporal samples are needed, but manual annotations are labor-consuming, and neglect the inherent temporal dynamics, impractical for continuous monitoring of large-scale geographical areas experiencing rapid surface alterations. Existing change detection methods depend on prior strong temporal pattern hypotheses, failing to achieve complex geographical temporal awareness. Moreover, the end-to-end manner with the strong temporal pattern hypotheses can lead to unreliable results. Instead, we argue that a more robust approach involves initially performing change detection on individual training samples to get dynamic samples across time series. By introducing dynamic samples as an interpretable and controllable intermediate variable, final classification results are ensured to be grounded in well-verified surface transition evidence. These dynamic samples can then be employed to train classifiers that yield improved dynamic mapping results. This two-stage approach provides greater control over intermediate variables, significantly enhancing the accuracy of dynamic change detection outcomes.
@@
Fundamentally, land surface change (e.g., deforestation, urban expansion, wetland degradation) can be understood as a form of spatiotemporal anomaly. As Fig. 1 shows, these changes are typically characterized by disruptions in temporal trends and simultaneous deviations across multiple spectral dimensions. For instance, the breakdown of expected correlations between bands (e.g., NIR and SWIR) or the disappearance of seasonal vegetation cycles in indices like NDVI often signal such changes. When focusing on representative training samples, recognizing these patterns as spatiotemporal anomalies and identifying these anomalies provide a foundation for detecting land surface dynamics in time-series remote sensing data to generate dynamic training samples. Consequently, we consider developing an automated framework capable of dynamically capturing heterogeneous Earth’s surface change patterns by extracting spatiotemporal anomalies of training samples. Based on the existing static training samples as the initialization, this framework then generates reliable dynamic samples for dynamic remote sensing geographic mapping. However, this idea faces two main challenges: First, remote sensing time-series dynamic patterns involve temporal patterns and spectral patterns. Independently learning the spectral and temporal embeddings can make feature fusion challenging, leading to the inconsistency of the learned spectral and temporal embeddings of remote sensing time-series data. Second, the lack of interpretable attribution in anomaly detection limits the model’s ability to precisely localize the source of changes. Without analyzing the contributions of different spectral bands and time points, anomalies may be detected without a clear causal explanation, reducing the robustness and interpretability of the model, especially in complex or heterogeneous environments.
@@
    Our proposed TasGen is the first method focusing on automatic dynamic sample generation for time-series dynamic remote sensing geographic mapping. By jointly modeling temporal and spectral dependencies via a hierarchical temporal-spectral Variational AutoEncoder (HTS-VAE) and a dual-dimension embedding, TasGen aligns and models patterns across spectral and temporal dimensions to enable accurate anomaly detection.
@@
    Two comprehensive evaluation experiments, including dynamic land cover mapping and wetland mapping across different six regions, demonstrate the effectiveness and robustness of our proposed TasGen, confirming TasGen’s potential to enhance automated geographical mapping practices and reduce the associated labor demands in remote sensing dynamic monitoring.
2 Preliminaries
@@ 2.1 Problem Definition
Figure 2: The anomaly representations in time-series remote sensing samples.
2.2 Temporal-spectral Patterns
In remote sensing time-series data, the temporal-spectral patterns of geographic type changes can be categorized into three distinct types: temporal anomalies, spectral anomalies, and temporal-spectral anomalies. As shown in Fig. 2 (a), temporal anomaly means a subset of spectral bands (e.g., Blue, NDVI) exhibit significant deviations from their historical daily or weekly patterns. For example, an abrupt drop in NIR and NDVI during the growing season may indicate drought or fire disturbance. Such anomalies often reflect system-level changes in land cover or environmental conditions. Fig. 2 (b) shows the spectral anomaly, which indicates that most bands follow their individual temporal trends, but the multivariate relationships among bands deviate from historical norms. For instance, a disrupted positive correlation between NIR and SWIR bands might occur due to soil moisture changes or crop stress. This pattern violation typically arises from localized spectral changes rather than global trends. As for the temporal-spectral anomaly, both temporal dynamics and spectral relationships are disrupted. While these anomalies are often easier to detect due to their combined effect, they require careful disentanglement of causal factors (e.g., land use conversion altering both seasonal greenness and surface temperature). Regardless of their type, anomalies in remote sensing time series provide critical indicators of environmental changes. Accurate detection and characterization of these anomalies are the basis for capturing the Earth’s surface changes.
3 TasGen
@@ 3.1 Overview of TasGen
Figure 3: The overall architecture of TasGen. (a) is the variational net, and (b) is the generative net, and (c) is the sample relabeling process. $X$ is the input sample with a shape of $C$ bands and $T$ time points. $\mathit{e_{t}}$ is the temporal embedding, and $\mathit{e_{s}}$ is the spectral embedding. $X^{\prime}$ represents the reconstructed time-series patterns of the dynamic sample. $\mathit{y}$ is the label of stable samples. $\mathcal{N}$ and $\mathcal{A}$ mean the normal patterns and the anomalies. $\mathit{y^{\prime}}$ is generated labels for the dynamic sample.
TasGen is a two-stage framework designed to automatically generate dynamic training samples for remote sensing dynamic geographic mapping. Given a static labeled sample at a specific time point, TasGen aims to reconstruct its temporal evolutions and generate the dynamic states across the full observation window, capturing underlying temporal-spectral patterns without additional manual annotation.
To this end, TasGen first learns disentangled spectral and temporal representations through a combination of hierarchical temporal-spectral Variational Autoencoder modeling and dual-dimension embedding. To enhance robustness against noisy or anomalous observations in the training data, such as those caused by clouds, sensor issues, or transient surface disturbances, we design a sample curation strategy that learns stable spectral embeddings before joint training. In the second stage, dynamic sample relabeling is conducted using a classifier informed by the learned embeddings, enabling temporal label propagation from a single static annotation. Additionally, we introduce a Gibbs-based anomaly attribution mechanism to interpret the contributions of spectral-temporal components and enhance model transparency. The overall framework is illustrated in Fig. 3.
@@ 3.2 Dual-dimension embedding
We aim to model the temporal-spectral patterns of remote sensing time-series dynamic samples to detect anomalies, contributing to generating dynamic remote sensing samples. Therefore, the core of our model is how to represent each remote sensing dynamic sample well by jointly modeling its low-dimensional spectral and temporal embedding dependencies. To explicitly model the dependencies, we propose a dual-dimension embedding architecture. This mechanism compresses the time-series input along both the temporal and spectral dimensions, generating disentangled temporal and spectral representations.
Figure 4: The sample data structure.
@@
Existing static annotations are supervisory signals for land monitoring. However, these annotations typically correspond to a single time point and may reflect only a transient or abnormal state of the surface, rather than the typical and temporally stable pattern of the land surface. This can introduce significant bias into the model, causing it to overfit to non-representative spectral-temporal patterns and degrade its generalization ability across time. Inspired by the scheduled sampling for sequence prediction and sequence reconstruction [Bengio et al., 2015, Li et al., 2021], note that here we use a sample curation strategy Formula 3 that mitigates the impact of such anomalies before training. Instead of directly learning the spectral embedding from the raw, potentially noisy input x, we first generate a reconstructed input $\textbf{e}^{\prime}_{t}$ using the temporal embedding $\textbf{e}_{t}$. This reconstruction is designed to preserve the stable temporal patterns while filtering out short-term fluctuations and anomalies. By learning spectral embeddings from this cleaned reconstruction, the model avoids encoding spurious correlations associated with transient disturbances in the annotations. To ensure the quality of the reconstruction at the early stages of training, we pretrain this reconstruction path using a temporal VAE, allowing $\textbf{e}^{\prime}_{t}$ to act as a strong proxy for the typical surface state. This reconstructed signal filters short-term noise (e.g., clouds, sensor anomalies) while preserving stable time patterns, significantly reducing the risk of overfitting to anomalous or biased labels, while still leveraging the existing static annotations for supervision.
We then extract the spectral embedding $e_{s}$ by applying a spectral-wise recurrent encoder (bidirectional GRU $G$ [Chung et al., 2014]) along the spectral dimension:
@@ 3.3 Hierarchical temporal-spectral Variational Autoencoder
While our dual-dimension embedding structure effectively disentangles temporal and spectral representations, these two views are learned independently. This separation facilitates noise-robust feature extraction and temporal-spectral alignment, but may result in a lack of contextual interaction between temporal and spectral embeddings. To address this, we introduce a hierarchical variational autoencoder (HTS-VAE) (Fig. 3 (a)(b)) that jointly models these two embeddings in a coherent latent space. Specifically, we treat the temporal embedding as a contextual prior for spectral embedding learning, thereby enabling temporal-spectral simultaneous learning while preserving the structural benefits of disentangled encoding.
@@ 3.4 Anomaly Detection Inference
After learning the joint temporal-spectral representations using HTS-VAE, we perform anomaly detection on dynamic remote sensing samples by evaluating reconstruction-based deviations. The core assumption is that observations conforming to the learned temporal-spectral normal patterns can be reconstructed well, while anomalies, such as abrupt land cover changes, noise, or atypical surface responses, will show high reconstruction errors.
Given an input window $x\in\mathbb{R}^{C\times T}$, we pass it through the trained HTS-VAE to obtain latent embeddings $e_{t}$ and $e_{s}$. These embeddings are decoded to generate a reconstruction $\hat{x}$. We define the anomaly score as the negative log-likelihood (or equivalently, the reconstruction probability):

### v2

@@ Abstract
Accurate remote sensing geographic mapping requires timely and representative samples. However, rapid land surface changes often render static samples obsolete within months, making manual sample updates labor-intensive and unsustainable. To address this challenge, we propose TasGen, a two-stage Temporal spectral-aware Automatic Sample Generation method for generating dynamic training samples from single-date static labels without human intervention. Land surface dynamics often manifest as anomalies in temporal-spectral sequences. To effectively capture these dynamics, TasGen first disentangles temporal and spectral features to isolate their individual contributions, and then couples them to model their synergistic interactions. In the first stage, we introduce a hierarchical temporal-spectral variational autoencoder (HTS-VAE) with a dual-dimension embedding to learn low-dimensional latent patterns of normal samples by first disentangling and then jointly embedding temporal and spectral information. This temporal-spectral embedding enables robust anomaly detection by identifying deviations from learned joint patterns. In the second stage, a classifier trained on stable samples relabels change points across time to generate dynamic samples. To not only detect but also explain surface dynamics, we further propose an anomaly interpretation method based on Gibbs sampling, which attributes changes to specific spectral-temporal dimensions. We evaluate TasGen on two dynamic geographic mapping tasks (i.e., land cover mapping with 5 classes and wetland mapping with 4 classes) across six diverse global locations, each covering an area of around $12,062\ km^{2}$ and representing different continents and climate zones. TasGen achieved an average mapping OA and Kappa of 91.64% and 0.9071, considerably outperforming other state-of-the-art algorithms with a maximum of 7.97% and 0.0762. Consistent superior accuracy across all regions demonstrates TasGen’s effectiveness in tracking both long-term trends and abrupt land surface changes, while significantly reducing the need for human intervention.
@@ 1 Introduction
The Earth’s surface is undergoing unprecedented rapid changes, with urbanization encroaching upon natural habitats at an annual rate of 1.2% [Seto et al., 2012], and global wetlands decreasing by approximately 0.5% per year since the 20th century [Nicholls, 2004, Yuan et al., 2025]. Timely and accurate geographic mapping serves as a fundamental basis for monitoring ecosystem health [Soubry et al., 2021, Mariano et al., 2018], optimizing urban planning [Gong et al., 2020, Yuan et al., 2024, 2020], protecting biodiversity [Wiens et al., 2009], and understanding climate change [Yang et al., 2013]. Remote sensing has become the primary way of acquiring dynamic surface information with its capability for periodic observations [Gong, 2012, Pekel et al., 2016]. For instance, the Landsat satellite series has provided over 40 years of continuous observations, offering essential data for global land change analysis [Wulder et al., 2019].
@@
Now there are three kinds of methods addressing this challenge. The first is manually annotating multi-temporal samples. It requires visual interpretation and manual annotations at each time point, which is constrained in small study areas and short time series easily. For example, Li et al. [2017] established the first all-season sample sets (FAST) for global land cover mapping, primarily based on manual interpretation of Landsat-8 imagery regarding high-resolution Google Earth images. Yang & Huang [2021] collected training samples by integrating land cover samples derived from China’s Land-Use/Cover Datasets (CLUDs) with visually interpreted samples based on satellite time-series data, Google Earth, and Google Maps imagery. The second is using stable samples across time series to train phase-general classifiers. In other words, these methods only need manual annotations once to conduct the time-series mapping. For example, Zhang & Weng [2016] developed a method for extracting training samples from spectrally stable areas, and used Google Earth imagery as reference data for land cover identification. Huang et al. [2020] developed a novel methodology to migrate global training samples by leveraging spectral similarity and spectral distance metrics, enabling the reuse of stable training samples for dynamic global land cover mapping across multiple periods. Even though stable samples only need one-time annotation, the scarcity of stable samples and the temporal identification of stable samples limit the scaling ability. The third is the change detection method modeling the time-series patterns. Zhu & Woodcock [2014] proposed the continuous change detection and classification (CCDC) algorithm, detecting changes through unsupervised time-series modeling and training classifiers relying on static reference data from stable periods. Kennedy et al. [2010] proposed LandTrendr, considering dynamic segmentation analysis of Landsat time series and training classifiers relying on trajectory fitting and seasonal-trend decomposition. He et al. [2024] employed temporal semantic segmentation to extract land cover change information from Sentinel-2 time series data. These methods directly identify change points and their types at the mapping level, modeling temporal dynamics in an end-to-end manner. While powerful, Zhu & Woodcock [2014] and Kennedy et al. [2010] struggle to adapt to serious Earth’s surface patterns because of the model initialization. Besides, they only considered the temporal dimension but neglected the synergistic effects of temporal-spectral patterns when land surface changes. He et al. [2024] has limited scalability because of the need for temporal-dense annotations. Additionally, such frameworks can benefit from the integration of intermediate constraints at the sample level, with better interpretation ability and enhanced control over the learning process.
@@
Due to the complexity of both spectral and temporal dynamics inherent in time-series data, obsolete samples cannot effectively represent surface characteristics at later time points, leading to significant degradation in mapping accuracy over time. High-quality and abundant dynamic and up-to-date samples are the key to accurate time-series dynamic mapping [Stanimirova et al., 2023]. Based on the observations above, dynamic samples are needed, but manual annotations are labor-consuming, and neglect the inherent temporal dynamics, impractical for continuous monitoring of large-scale geographical areas experiencing rapid surface alterations (Fig. 1 (b)). Existing change detection methods often rely on strong assumptions about temporal patterns or temporal-dense annotations, and neglect the synergistic effects of temporal-spectral patterns. This limitation is particularly problematic for wetlands, which exhibit recurrent phenological cycles, and subtle intra-annual transitions that are only distinguishable through the joint modeling of spectral signatures and their temporal evolution [Yuan et al., 2025] (Fig. 1 (c)). Besides, while end-to-end frameworks offer efficiency, introducing intermediate sample-level constraints allows for greater flexibility and interpretability, improving robustness in diverse and evolving environments. Therefore, we argue that a more robust approach involves initially performing change detection on individual training samples to get dynamic samples across time series. By introducing dynamic samples as an interpretable and controllable intermediate variable, final classification results are ensured to be grounded in well-verified surface transition evidence. These dynamic samples can then be employed to train classifiers that yield improved dynamic mapping results (Fig. 1 (a)). This two-stage approach provides greater control over intermediate variables, significantly enhancing the accuracy of dynamic change detection outcomes.
Fundamentally, land surface change (e.g., deforestation, urban expansion, wetland degradation) can be understood as a form of spatiotemporal anomaly. As Fig. 2 shows, these changes are typically characterized by disruptions in temporal trends and simultaneous deviations across multiple spectral dimensions. For instance, the breakdown of expected correlations between bands (e.g., NIR and SWIR) or the disappearance of seasonal vegetation cycles in indices like NDVI often signal such changes. Recognizing these patterns as spatiotemporal anomalies provides a conceptual foundation for detecting land surface dynamics in time-series remote sensing data. Therefore, a key step towards robust dynamic geographic mapping is to generate reliable dynamic training samples from limited static labels by identifying such anomalies. However, generating dynamic samples from static supervision faces two main challenges. First, remote sensing time-series dynamics are both multivariate and unified, which means they involve distinct temporal and spectral components that reflect different aspects of surface change, yet jointly characterize a coherent land surface state. Purely decoupled learning of temporal and spectral embeddings may fail to capture their integrated semantic meaning, while purely coupled modeling risks entangling fundamentally different variation sources. This tension makes it challenging to achieve consistent and effective temporal-spectral representation learning. Second, the lack of interpretable attribution in anomaly detection limits the model’s ability to precisely localize the source of changes. Without analyzing the contributions of different spectral bands and time points, anomalies may be detected without a clear causal explanation, reducing model robustness and interpretability, especially in complex or heterogeneous environments.
Figure 1: The workflow of each dynamic mapping paradigm. (a) is the workflow of our proposed method; (b) is the workflow of the manual annotation; (c) is the workflow of the change detection methods. We introduce dynamic samples as an interpretable and controllable intermediate variable for final classification results.
@@
    Our proposed TasGen is the first method focusing on automatic dynamic sample generation for time-series dynamic remote sensing mapping. By first decoupling temporal and spectral dependencies and then jointly modeling them via a dual-dimension embedding and a hierarchical temporal-spectral Variational AutoEncoder (HTS-VAE), TasGen aligns and models patterns across spectral and temporal dimensions to enable accurate anomaly detection.
@@
    Two comprehensive evaluation experiments, including dynamic land cover mapping and wetland mapping across six different regions, demonstrate the effectiveness and robustness of our proposed TasGen, confirming TasGen’s potential to enhance automated geographical mapping practices and reduce the associated labor demands in remote sensing dynamic monitoring.
2 TasGen
@@ 2.2 Overview of TasGen
TasGen is a two-stage framework designed to automatically generate dynamic training samples for remote sensing dynamic geographic mapping. The overall framework is illustrated in Fig. 3. Given a static labeled sample at a specific time point, TasGen aims to reconstruct its temporal evolutions and generate the dynamic states across the full observation window, capturing underlying temporal-spectral patterns without additional manual annotation.
[…]
To this end, TasGen first decouples the temporal and spectral dependencies through a dual-dimension embedding, and then learns disentangled spectral and temporal representations jointly through a hierarchical temporal-spectral autoencoder (HTS-VAE). Anomalies are detected based on the reconstructed temporal-spectral patterns. In the second stage, dynamic sample relabeling is conducted using a classifier trained by the learned embeddings, enabling temporal label propagation from a single static annotation. Additionally, we introduce a Gibbs-based anomaly attribution mechanism to interpret the contributions of spectral-temporal components and enhance model transparency.
@@ 2.3 Dual-dimension embedding
We aim to model the temporal-spectral patterns of remote sensing time-series dynamic samples to detect anomalies, contributing to generating dynamic remote sensing samples. To explicitly model the dependencies, we first propose a dual-dimension embedding architecture, which disentangles temporal and spectral representations by compressing the time-series input along both the temporal and spectral dimensions.
@@
Static annotations often reflect transient surface states, risking model overfitting to anomalous spectral-temporal patterns. To address this, we adopt a sample curation strategy: instead of directly learning from raw inputs, we first reconstruct a stable time series $\textbf{e}^{\prime}_{t}$ using the temporal embedding $\textbf{e}_{t}$, then derive spectral embeddings from this denoised signal. This approach preserves typical surface dynamics while filtering short-term noise (e.g., clouds), thereby reducing annotation bias. We further pretrain the reconstruction path using a temporal VAE to ensure early-stage stability and improve generalization. We then extract the spectral embedding $e_{s}$ by applying a spectral-wise recurrent encoder (bidirectional GRU $G$ [Chung et al., 2014]) along the spectral dimension:
@@ 2.4 Hierarchical temporal-spectral variational autoencoder
While our dual-dimension embedding structure effectively disentangles temporal and spectral representations, these two embeddings are learned independently. This separation may result in a lack of contextual interaction between temporal and spectral embeddings. To address this, we introduce a hierarchical variational autoencoder (HTS-VAE) (Fig. 3 (a)(b)) that jointly models these two embeddings in a coherent latent space. Specifically, we treat the temporal embedding as a contextual prior for spectral embedding learning, thereby enabling temporal-spectral simultaneous learning while preserving the structural benefits of disentangled encoding.
@@ 2.5 Anomaly detection inference
After learning the joint temporal-spectral representations using HTS-VAE, we perform anomaly detection on dynamic remote sensing samples by evaluating reconstruction-based deviations. Given an input window $x\in\mathbb{R}^{C\times T}$, we pass it through the trained HTS-VAE to obtain latent embeddings $e_{t}$ and $e_{s}$. These embeddings are decoded to generate a reconstruction $\hat{x}$. We define the anomaly score as the negative log-likelihood (or equivalently, the reconstruction probability):